a galvanometre has a resistance of G ohm and range of volt.calculate the resistance to be used in series with it to extend its range to v volt.
ans= G(n-1).
please explain in detail
 

Grade:12

1 Answers

Susmita
5 years ago
Hey  your question may have some error in it.
If a galvanometer initially has a range of 0 to 1 volt with internal resistance G and if you want to increase the range to 0 to v volt,then this is the answer.
You know that V=IR
Or,I=V/R
Current I will remain fixed.Only V,R will change in this problem.
Initially V=1 volt and R=G.So
I=1/G.
Next V=v volt and R=G+r,where r is the added resistance.So
I=v/(G+r)
Equate the two equations.
1/G=v/(G+r)
Or,vG=G+r
Or,r=vG-G
Or,r=G(v-1)
Please approve the answer if you are helped.Feel free to ask if there is any further query.